GOHSEP Daily Brief – September 12, 2021

GOHSEP Daily Brief

September 12, 2021

As of 5:00 p.m.

State EOC Activation Level: Level I (Full Activation)

WebEOC Incident: 21-018 TC Ida – Statewide – August

State Declaration Proclamation Number: 165 JBE 2021


SITUATION:  Hurricane Ida made landfall along the southeastern coast of Louisiana near Port Fourchon around 11:55 am with estimated maximum sustained winds of 150 mph. Areas of southeast Louisiana are still under flood advisories. Additional rainfall is possible over southeast Louisiana as well heat advisories. Hurricane Ida caused devastating destruction across southeast Louisiana. Power outages are wide spread and several parishes are having continuing communication issues. Damage is still being assessed. Local, state, and federal partners are responding to the needs of the parishes.





·         WebEOC Incident 21-018 TC Ida – Statewide – August created 8/16/2021.

·         Level III Crisis Action Team Activation 8/26/21.

·         Level I Full Activation 8/27/21.


ESF-02 Communications

·         LWIN System Status at 5:00 PM on 09-12-21

o   Sites in wide area operation: 141/144

§  Bourg: Comm Failure

§  Houma: Site Trunking

§  Greensburg: Site Trunking

o   Sites are in wide area operations, but are on generator back up power:

§  Geismar, St. James, Northern Simulcast Gretna, NOLA East & Galleria, Laffite
Bridge City, Larose, Gray, Houma, and Montegut

·         Tactical Equipment:

o   Emergency tower trailers

§  Located in Baton Rouge: Box Trailer 99

§  Located in Alexandria: None

·         Deployed: 4

o   Trailer 96 (LSA) deployed to Houma

o   Trailer 97 in use at Calcasieu Sheriff’s Office before Ida

o   Trailer 100 deployed at Galliano

o   Trailer 98 deployed at Grand Isle on standby to be redeployed

o   Generators on Wheels (GOW)

§  Located in Baton Rouge: 2

·         Deployed: 2

o   Deployed at Larose

o   Deployed at Jackson

§  Located in Alexandria: 2

·         Deployed at HQ

o   Flyaway kits

§  Located in Baton Rouge: 1

·         Deployed: 3

o   Larose (Standby)

o   Theriot

o   Laplace

§  Located in Alexandria: None

·         Blue Mobile Command Post located in Baton Rouge

·         White Mobile Command Post deployed to Houma Civic Center


ESF-04 Firefighting Urban

·         444 personnel in 92 fire apparatus deployed across 7 Parishes.

·         In previous 48 hours, Fire Crews responded to 622 emergency calls, 12 were Structure fires

·         OH IMT has replaced FL and are now on duty.

·         SFM in coordination with the Jefferson Parish to assist them to obtaining an IMT for restoration efforts.

·         SFM working with GOHSEP to obtain damage assessments for all fire departments in the effected areas.

·         CO alarm Distribution and generator safety messaging continues

·         Blue tarps and bottled water being distributed to older residents.

ESF-06 American Red Cross

·         ARC currently operating 14 shelters in 8 different parishes housing over 1032 clients.

·         Distributed critical supplies on numerous routes. (water, heater meals & Comfort kits), serving over 2,701 households on yesterday and 16,397households to date.

·         Operating 3 large partner kitchens with one more set to go online. Supplying hot meals in multiple Parishes; ARC has served over 34,393 yesterday and 304,282meals to date.

·         Ongoing and continuous Damage Assessment in all the affected areas, and have assessed over 32.7 K residences in 19 Parishes

ESF-06 DCFS Mass Care

·         Alexandria Mega Shelter, Bastrop Shelter and Jewella Shelters are open as Critical Transportation Needs Shelters. Mega Shelter MSN, Bossier Civic Center MSN, LSU Field House MSN and MCC MSN are open as Medical Special Needs Shelters.

·         Coordinating with Orleans Parish on re-entry from CTN shelters. Continue to move forward with shelter consolidation.

o   Shelter Population Update 2021-09-12 0800 hours

§  21 Total Shelters(s) – 1528

·         Population

o   3 State – 324

o   4 MSN – 193

o   0 Parish –

o   14 ARC – 1011

o   0 Faith –

o   0 State NCS –

o   0 ARC NCS –

·         Shelter Reporting cycle is 0800, 1800, and 2400.

·         Shelter Report is posted in WebEOC on File Library under ESF-6 DCFS.

·         DCFS submitted request for DSNAP to FNS on 9/9/21, awaiting approval. Once approved, DSNAP schedule will be communicated to the IA parishes.

ESF-06 LWC Mass Feeding

·         LWC mass feeding operations continue at 6 shelter facilities with daily meal count:

o   Alexandria Mega Shelter – 700

o   Bastrop Shelter – 200

o   Jewella Shelter – 350

o   Bossier Civic Center MSNS – 130

o   LSU Field House MSNS – 165

o   Morial Convention Cntr FMS – 250

·         Feeding 5,385 meals today.

ESF-08 LDH Public Health/Medical

·         26 Verified Storm Related Deaths (as of 09/12/2021 @ 1500)

·         Surge Ambulances (as of 1500 on 09.12.2021)


§  State Surge

·         Staged: 0

·         Assigned: 22

§  Federal

·         Staged: 26

·         Assigned: 198

§  Paratransit

·         Staged: 24

·         Assigned: 19

·         Healthcare Facilities (As of 1500 on 09.12.2021)

o   Hospitals as of 09.11.2021 @ 1500

§  Hospital Evacuations: 15 Facilities

§  Total Hospitals on generators: 11

o   Nursing Homes as of 09.12.2021 @ 1500

§  Nursing Home Evacuations: 10 Facilities

§  Total Nursing Homes on Generators: 10

o   Assisted Living as of 09.12.2021 @1500

§  Assisted Living Evacuations: 23 Facilities

§  Total ARCPs on Generators: 8

o   Intermediate Care Facilities as of 09.12.2021 @1500

§  93 ICF Evacuated

·         Water Systems/Engineering (as of 1500 on 09.12.2021)

o   19 storm-related Water Outages There are 7 community water systems (CWS) with a water outage. There are at least 8 non-community water systems that are Closed due to no power.

o   86 storm-related Boil Water Advisories

o   221 water systems have been CLEARED from the storm-related boil water advisory



·         Surveillance Staff continue to perform field assessments of facilities, waste water treatment plants, Emergency Debris Sites, and Landfills.

·         Surveillance and Emergency Response Staff continue to perform field assessments of oil spills with LOSCO. Sites requiring active clean-up operations are being closely monitored.

·         The Waste Permits Division is working with Parishes to review, approve, and activate Emergency Debris Sites.

·         The Mobile Air Monitoring Labs are located in Norco

·         LDEQ remains in contact with State and Federal ESF-10 partners concerning post storm activities.

·         The Department remains in contact with the Nuclear Power Plants:

·         Waterford – The facility started supplying power to the grid on 9/7/21, anticipates being at 100% power today, currently supplying 48% to the grid.

·         Mission Assignment 1 – EPA ASPECT Aircraft -The EPA ASPECT aircraft completed the primary task of the Mission Assignment to fly the industrial corridor. A secondary task of this MA was conducted yesterday to collect aerial imagery of oil spills in the region. The plane will remain on stand-by in Addison, TX for future flights if requested.

·         Mission Assignment 2 – EPA Air Monitoring/Sampling Assistance – EPA contractors are performing air monitoring in St. John and St. Charles Parishes. Air samples in these areas will begin being collected for analysis today.


ESF-11 Department of Agriculture and Forestry

·         LDAF continues to operate an Emergency Fuel Depot 24-hours daily in Hammond and manages emergency fuel for life-saving and support missions.

·         LDAF continues to provide supplies to the parishes of Jefferson, St. Charles, St. James, Assumption and Terrebonne.

·         The Louisiana Farm Bureau Federation, Louisiana Cattlemen’s Association, LSU AgCenter, and LSART continue to assess damage for and respond to the agricultural and livestock communities. The LDAF Hammond complex is a staging and distribution site for these assets as
well as needed livestock supplies.

·         LDAF continues operations of three CTN co-located pet shelters: Alexandria Mega-Shelter, Haughton Shelter, and Bastrop MAC. DCI is sheltering for medical special needs pets. Today, we are preparing for the departure of animals from the Haughton pet shelter. Pet Census: Alexandria: 15 dogs/5 cats; Haughton: 10 dogs/12 cats, 1 dog at animal control; Bastrop: 12 dogs/5 cats; DCI: 1 dog

·         Louisiana State Animal Response Team (LSART) conducted veterinary medical vaccination clinics at the state-run and parish shelters when requested and continues to provide technical assistance and large animal assessments.

·         LSART, IFAW, and ASPCA are supporting ongoing parish pet operations including small animal assessments.

·         Hill’s Pet Nutrition continues its support to parishes with pet food and supplies.

·         LDAF is supporting PODS as requested by LANG.

·         LDAF Food Branch is monitoring DHD, assisting food banks and the Council on Aging with food storage.

·         LDAF Livestock Brand Commission is assisting producers with transporting stranded cattle in Plaquemines Parish.


·         19 storm-related Water Outages for an Estimated affected population ~19,724. There are 7 community water systems (CWS) with a water outage. There are at least 8 non-community water systems that are Closed due to no power.

·         86 storm-related Boil Water Advisories for an Estimated affected population of ~318,654.

·         221 water systems have been CLEARED from the storm-related boil water advisory for an Estimated affected population of ~829,755.

·         See attached Water Outage and Boil Water Advisory Reports in WebEOC for a complete list of affected water systems. Water systems with on a water outage will be under a boil water advisory once system pressure is restored.

ESF-12 Public Service Commission

·         Power outages statewide as of 09:00 on 09/12/2021 were: 140,189

·         Outage map: https://lpsc.louisiana.gov/Outages/Map/g7h


ESF-13 Department of Corrections

·         P&P EOC update for missions worked and scheduled on 9/12/2021 as of 0600hrs.

o   P&P HQ staff manning the P&P Operations Center during business hours and remotely monitoring all other hours for 24hr/7 day coverage.

o   8 P&P Officers (4 officers per shift for 24 hours) assisting LSP EOC with security at State Museums in the NOLA French Quarter. Mission began on 09/1/2021 and is ongoing.

o   10 P&P Officers (5 per shift for 24 hours) assisting LSP with the Red Cross Shelter at the River Center in Baton Rouge. Mission began on 9/3/2021 and is ongoing.

o   4 P&P Officers (2 per shift for 24 hours) supporting Albany PD with community policing. Mission began on 9/3/2021 and is ongoing.

o   4 P&P Officers (2 per shifts for 24 hours) providing security support at the sex offender shelter in Homer. Mission began on 8/30/2021 and is ongoing.

o   4 P&P Officers (2 per shift for 24 hours) assisting LSP with security at Jewella Shelter in Shreveport. Mission began on 9/7/2021 and is ongoing.

·         DOC IMC UPDATES 9/12/21

o   Parish Facilities That Remain Evacuated: Terrebonne

o   Parishes housing evacuated Offenders: Madison, Catahoula, LaSalle, Beauregard, Assumption

o   Transport Missions: NONE SCHEDULED

o   Other impacted parish reports / movements: NOTHING NEW REPORTED

o   DOC FACILITY REPORTS: DWCC SO Shelter reporting 1 evacuee

ESF-13 Louisiana State Police

·         LSP EOC will return to normal operations on 9/13. Please contact ESF-Leads for any afterhours requests.

·         LSP distributed important targeted messaging from CISA to our field personnel regarding not cutting communications lines during recovery efforts.

·         3 escorts completed last night from the Roseland Regional Staging Area.

·         LSP continues proactive enforcement across all affected areas with multiple task forces, including assistance from several out of state police agencies.

·         Alabama Highway Patrol arrived today and will assist with Hurricane Ida security operations. Mississippi Highway Patrol will arrive on 9/13 with additional assets.

·         Shelter security operations continue at 9 sites statewide (Johnson and Marbles Community Centers closed at 12 PM today).

·         Since 8/25, LSP has generated numerous storm-related social media posts reaching over 12 million social media users.

ESF-16 Military Support

·         8246 Total Forces supporting current operations.

o   HURRICANE IDA – 7429:

§  SAD: 3045

§  Other Support: 1156

§  T-10 DSC: 1

§  COVID PAX Supporting: 273

§  T10 Support: 412

§  EMAC Forces: 2542

·         183 Total on-going missions:

o   138 Hurricane Ida Missions

o   44 missions LANG COVID-19

o   1 mission EMAC (California Wildfire)

·         Commodity Distribution:

o   Currently have 50 PODs and 4 HUBs deployed in 10 parishes

o   Closed 1 HUB in Livingston

o   30 resupply trucks overnight

o   3 x Direct Deliveries to 3 parishes in last 24hr

o   Project closing 5 PODs today

o   Commodities distributed:

·         Bulk Water:

o   37 water buffalos, 10 hippos, 5 camels in 12 parishes (St. James, Terrebonne, St. Charles, Lafourche, Jefferson, Plaquemines, Orleans, St. John, Ascension, St. Tammany, Washington, Tangipahoa)

o   76,400 gal (+15,700) to 6 sites – 1 in Plaquemines and 5 in Terrebonne

·         Task Force Generator:

o   680 requested

§  297 on mission

§  3 installations in progress

§  0 pending assessment

§  0 on hold

§  90 reset

§  290 cancelled

o   Increased 17 generators since last report

o   4 generator installations planned today

·         Engineer Operations:

o   Route Clearance: 5302 (+6) mi assessed and 2,936 mi of roads cleared in 20 Parishes. 93 (+2)

o   Approximately 98% of requested route clearance complete

o   Working clearing operations in 2 State Parks in Livingston and St. Tammany

o   Continue debris removal in drainage canals in Tangipahoa and St. Charles

·         Security Operations:

o   877 Guardsmen conducting security missions in 17 Parishes

§  Orleans Parish – 214 PAX ISO NOPD

§  Jefferson Parish – 56 PAX ISO law enforcement

§  Plaquemines – 25 PAX TCP operation

§  Tangipahoa – 42 PAX presence patrols

§  Lafourche – 154 PAX ISO LE

§  Terrebonne – 90 PAX ISO LE

§  St. Helena – 90 PAX ISO LE

·         Medical Support Operations: 65 PAX

o   Medical Team supporting MSNS @ MCC, Senior Military Advisor Supporting Incident Commander @ MCC-MSNS

o   CERFP MED-L team ISO Terrebonne Gen Hospital

o   1 x Provider & 2 x Medics to NSU Student Union

·         Government Support Operations:

o   9 teams, 1 closed Tangipahoa yesterday ( St. Helena, Livingston, Orleans, Jefferson, Plaquemines, St. Charles, St. John, Lafourche, Terrebonne)

o   9 x Senior Military Advisors deployed to (Tangipahoa, St. Helena, Orleans, Jefferson, Plaquemines, St. Charles, St. John, Lafourche, Terrebonne)

o   12 x Parish LNO Teams

·         FORCE FLOW:

o   EMAC: 2542 PAX from 15 states

§  Anticipate Phase 2 EMAC forces begin arriving today

·         Combat Sustainment Support Battalion (-) 173 personnel (OH)

§  Releasing 1 x CH47 to return to South Carolina

·         T10: 481 (Ida – 412) (COVID – 69)

o   Conducting Engineer Operations and logistics by TF TRUCK

·         COVID Update:

o   Operating 9 COVID test sites

o   Operating 1 COVID vaccine sites

o   Supporting­ 0 food banks

Cemetery Task Force

·         Two Ascension cemeteries contained 4 destroyed vaults with two exposed caskets.

·         As of today, the total compromised graves documented from this event is 107 and the total cemeteries surveyed is 33.


Assumption Parish

·         Assumption Parish lifted the Mandatory Evacuation for all of Assumption Parish.


Lafourche Parish

·         PORT WATERWAYS: The US Coast Guard has set Port Fourchon as Port Condition: Normal with Restrictions. Both Belle Pass and Port Fourchon waterways are open for 24/7 operations. CLICK HERE to view Marine Safety Information Bulletin MSIB 033-21 for details

·         WATER: The Lafourche Parish Water District continues to address issues in getting water back online, and they are making great progress.

o   Expect running water in Port Fourchon early this week.

o   As water service comes back online, Lafourche Parish Water District customers are asked to conserve water and refrain from non-essential use until service is fully restored. Please report any possible leaks to the Water District at 985-532-6924.

·         CURFEWS: There is a curfew in place from 7:00 p.m. to 6:00 a.m. in Lafourche Parish.

o   There is currently a dusk to dawn curfew in effect for both vehicular and marine traffic in Port Fourchon, but facilities may contact Port Fourchon Harbor Police at 985-396-2750 or harborpolice@portfourchon.com to request exceptions for nighttime deliveries.


·         ROADWAYS: Roadways are clear to and within Port Fourchon.

o   Please use caution when traveling recently cleared roadways. LA 1 to Grand Isle has limited access for property and business owners and emergency services.

·         TENANT FACILITIES: Port tenants have been given full access to return to their facilities, and several of our larger tenants have either returned to operations or are readying themselves for operations in the port. Tenants’ employees should contact their employers for specific information on returning to work in Port Fourchon.

·         BAYOU LAFOURCHE: Bayou Lafourche remains closed north of Leeville per US Coast Guard.

o   This portion of Bayou Lafourche north of the port is closed to marine traffic due to multiple obstructions in the channel, especially between the Ted Gisclair Floodgate in Larose and the Leon Theriot Lock in Golden Meadow.

o   NOAA Coast Survey is finalizing the survey process in Bayou Lafourche but is working with limited access due to shoaling and obstructions throughout the channel from Leeville to Larose. We are awaiting survey findings and will send out an update when that information is received.

·         POWER: There is no power, and Entergy’s estimated restoration date for Lafourche Parish is September 29.

·         COMMUNICATIONS: Most cell service has been restored, and outages are becoming less frequent.

·         COMMERCIAL FUEL: Stone Fuel is operational. C-Port Stone is also operational. Martin Energy will be operational by the beginning of next week.

·         FOOD, GAS, & SUPPLIES: There are no convenience stores, gas stations, or restaurants open at this time.

·         CLEANUP & DEBRIS: The Port Commission asks that as you begin to clean the debris from your business property, that you NOT place debris along the side of the roadway. You must stack debris within the footprint of your facility, as there is no debris staging area in port.

·         EMERGENCY SERVICES: Temporary fire and emergency services have been established in Port Fourchon, and Port Fourchon Harbor Police are continually working security for the area from a law enforcement perspective.